SUBMILLIMETER OBSERVATION OF HCO^{+} AND DCO^{+} IN THE EXCITED VIBRATIONAL STATES

T. HIRAO, Department of Chemistry, Okayama University, Okayama, Japan 700-8530; \underline{SHANSHAN YU}, and T. AMANO, Department of Chemistry, University of Waterloo, Waterloo, ON, Canada N2L 3G1.

Hirota and Endo reported observations of the rotational transitions of HCO^{+} and DCO^{+} in excited vibrational states.^{a} The lowest three rotational transitions of (02^{2}0) could not be detected in their experiment. They ascribed this nonobservation to the Stark broadening caused by the electric field in a hollow cathode discharge. More recently Dore and coworkers^{b} also did not see the ℓ = 2 lines in an extended negative glow discharge and interpreted this result in terms of the Stark effect. However, symmetric top ions such as CH_{3}CNH^{+} ^{c} and SD_{3}^{+} ^{d} were observed with no difficulty. Also no anomalies were observed for similar lines for HCN ^{e} and HNC ^{f} produced in an extended negative glow discharge. In the present investigation, we extended the measurements up to 800 GHz. The HCO^{+} and DCO^{+} ions were produced in an extended negative glow discharge in a gas mixture of H_{2} or D_{2} and CO ( a couple of mTorr each ) in Ar buffer ( 12 mTorr ). The measurements were done mostly at liquid nitrogen temperature. For HCO^{+}, the rotational lines in the excited vibrational levels up to (040) and (002) which are located at about 4300 cm^{−1} above the ground state have been measured. The measurements for DCO^{+} are not as extensive so far. Our observations confirmed that (02^{2}0) lines and, in general, larger ℓ lines in the (030) and (040) states were weaker than expected. However, a most notable result obtained in the present investigation is that some low-J lines of (02^{2}0) have been detected as induced emission for both HCO^{+} and DCO^{+}. This observation clearly leads to a conclusion that the previous non-observation of low-J lines in (02^{2}0) is not due to the Stark effect, but due to specificity of the reaction mechanism and subsequent collisional relaxation processes.

aE. Hirota and Y. Endo, J. Mol. Spectrosc. 127, 527 (1988).

bL. Dore, S. Beninati, C. Puzzarini, and G. Cazzoli, J. Chem. Phys. 118, 7857 (2003).

cT. Amano, K. Hashimoto, and T. Hirao, J. Mol. Structure, 795, 190 (2006)

dF. Tinti et al, J. Mol. Spectrosc. 240, 202 (2006)

eZ. Zelinger et al, J. Mol. Spectrosc. 220, 223 (2003)

fT. Amano and Z. Zelinger, J. Mol. Spectrosc. 211, 273 (2002)